Arlington Clothing, Inc., shows the following information for its two divisions for year 1:

   

Lake Region Coastal Region
Sales revenue $ 4,080,000 $ 13,020,000
Cost of sales 2,615,000 6,490,000
Allocated corporate overhead 244,800 781,200
Other general and administration 546,900 3,780,000

The results for year 2 have just been posted:

Lake Region Coastal Region
Sales revenue $ 4,080,000 $ 9,780,000
Cost of sales 2,615,000 4,820,000
Allocated corporate overhead 306,000 733,500
Other general and administration 535,000 3,780,000


Required:

a. Compute divisional operating income for the two divisions. (Enter your answers in thousands of dollars rounded to 1 decimal place.)

   

b-1. What are the gross margin and operating margin percentages for both divisions? (Enter your answers as a percentage rounded to 2 decimal places (i.e., 32.12).)

b-2. How well have these divisions performed? (You may select more than one answer. Single click the box with the question mark to produce a check mark for a correct answer and double click the box with the question mark to empty the box for a wrong answer.)

The reported divisional income for the Lake Region went down because sales fell in Coastal Region.
Corporate overhead is allocated on the basis of relative revenues.
Corporate overhead is allocated on the basis of absolute revenues.
The performance of the Lake Region is affected by the results in the Coastal Region.
The operating margin is greater in Lake Region.
The gross margin percentage is higher in Coastal Region.
Answer #1

a) Compute divisional operating income for the two divisions

For year 1

Lake Region Costal region
Sales 4,080,000 13,020,000
(less) Cost of sales 2,615,000 6,490,000
Gross margin 1,465,000 6,530,000
(less) Allocated corporate overhead 244,800 781,200
(less)Other general and Administration 546,900 3,780,000
Operating income/margin 673300 1,968,800

For year 2

Lake Region Costal region
Sales 4,080,000 9780,000
(less) Cost of sales 2,615,000 4,820,000
Gross margin 1,465,000 4,960,000
(less) Allocated corporate overhead 306,000 733,500
(less)Other general and Administration 535,000 3,780,000
Operating income/margin 624,000 446,500

b-1. What are the gross margin and operating margin percentages for both divisions?

For year 1

Lake Region

Gross Margin percentage = Gross margin / Sales

= 1,465,000/4,080,000

= 0.359

Costal Region = 6,530,000/13,020,000

= 0.5015

For year 2

Lake Region

Gross Margin percentage = Gross margin / Sales

= 1,465,000/4,080,000

= 0.359

Costal Region = 4960,000/9780,000

= 0.5071

For year 1

Lake region

Operating margin = Operating income/sales

= 673300/4,080,000

=0.1650

Costal Region = 1,968,800/13,020,000

= 0.1512

For year 2

Lake region

Operating margin = Operating income/sales

= 624,000/4,080,000

=0.1529

Costal Region = 446500/9780,000

= 0.0456

b-2. How well have these divisions performed?

Correct statements:
>Divisional income is greater in Coastal region
>Gross Margin percentage is higher in Coastal region.
>Corporate Overheads appear to be allocated on the basis of revenue